Recently Ryan Smith, the plant manager of the manufacturing division of Waterways Corporation, has been focusing on changes to overhead costs. He realizes that Ben Clark's new designs call for more automation in the plant, but he is also investigating if there are any opportunities for cost savings.

Ryan thought it might be helpful to his cost-cutting measures if he could predict what manufacturing overhead would be in the following months. But first he needed to determine the appropriate activity base. He thought there could be two possibilities: direct labour or the number of hours of operation.

From historical data, he retrieved the following information:

Direct Labour

Hours of Operation

Manufacturing Overhead

January

$25,000

500

$145,000

February

  24,000

520

  148,000

March

  30,000

700

  170,000

April

  32,000

690

  176,000

May

  27,000

575

  150,000

June

  25,000

550

  140,000

Ryan then asked CFO Jordan Leigh for information available to determine the cost of goods manufactured. Ryan was provided with the following information.

  • 1.The balances in the applicable inventory accounts at the beginning of the month were: Raw materials inventory $35,000; Work in process inventory $52,000.
  • 2.Raw material purchases for the month were $191,000.
  • 3.Of the raw materials used in production, 75% could be traced to the actual production, and the rest was indirect materials.
  • 4.Ending raw materials inventory was $50,000.
  • 5.Actual costs for wages and salaries were $70,000, of which 60% was considered overhead; the balance was direct labour.
  • 6.Hours of operation for the month were 600.
  • 7.Total manufacturing costs for the month were $315,000.
  • 8.Costs transferred into finished goods inventory for the month were $325,000.

Questions:

1. ). a.  

Using the high-low method, and based on the historical data provided, determine two possible cost formulas for manufacturing overhead.

b.  

Using the cost formulas developed in part (a), determine which activity base would be better for predicting manufacturing overhead.

2.)

Prepare a condensed cost of goods manufactured schedule.
Answer #1

Appropriate activity base for cost-cutting measues will be Hours of operations based on historical data given in the question.

For opportunities of cost saving, Ryan should consider hours of operations as the activity base to investigate cost-cutting measures. Highest activity is 700 hours of operations at $170000 manufacturing overhead whereas lowest activity is 500 hours of operations at $145000

Calculation of variable cost per unit using the high- low method

Variable Manufacturing overhead= (Manufacturing overhead of highest activity- Manufacturing overhead of lowest activit) / (Highest activity unit- Lowest activity unit) i.e. $(170000-$145000) / (700-500) equals to 25000/200. Hence Variable manufacturing overhead = $125 per unit

Solving for the fixed manufacturing overhead for highest activity i.e. Manufacturing overhead = Variable manufacturing overhead + Fixed Manufacturing overhead i.e. 170000= (125*700) + Fixed Manufacturing overhead

Fixed Manufacturing overhead = $170000- $87500 i.e. $82500. Now we have determined fixed manufacturing overhead, we can investigate and take cost-saving measures, hence possible cost formula for calculating Manufacturing overhead can be

Let Direct labor costs be x and hours of operation be y then

Manufacturing overhead= x + 125y+ (82500-x) total fixed manufacturing overhead will remain $82500 in this case

Current month calculation

Hours of operation is 600

Direct labor costs $28000 (40% of 70000) salary overhead= 42000 (60% of 70000) Total manufacturing costs for month i.e. $ 315000

Calculation of Direct Material: Opening direct material + Purchases - closing direct material

i.e. 35000+ 191000 - 56500( Indirect Material) - 50000 = $119500 (Direct Materials for current month)

Total Manufacturing costs = Direct Labor + Direct Material + hours of operation* Variable cost + salary overheads+  Manufacturing overheads

315000= 28000+119500+ (600*125) + 42000 + Manufacturing overheads

Manufacturing overheads= 315000-28000-119500-75000-42000 i.e. $125500

Using the cost formula, we should use hours of operations activity base to determine manufacturing overhead and result in cost-cutting measures.

Costs of goods manufactured schedule

Opening raw material inventory = $35000 + Purchases= $191000 - Indirect Material = $56500 - Closing material balance = $50000 i.e. 119500

Add Direct Labor Cost = $ 28000

Add Salary Overhead = $42000

Variable manufacturing overheads = 600*125 i.e. $75000

Manufacturing overhead =$125500

Total costs of goods manufactured =$ 315000